UT Feb 16 2022 - Split Night LBTI (1st half) - DEC (2nd half)



Staff:

  • OSA: S.Allanson
  • ISp: C.Bolyard
  • Mtn Manager: B.Bingham
  • Observer/ISA: J.Power
  • AO: Dr. Doug
  • Special Guests: B.Rothberg, R.T. Gatto

Night Info

  • Sunset 18:05
  • Nautical Twilight Ends 18:59
  • Astro Twilight Ends 19:28
  • Night Center 00:33 - HANDOVER FROM LBTI TO LUCI-AO (or Bino-LUCI) 00:27
  • Astro Twilight Begins 05:39
  • Nautical Twilight Begins 06:08
  • Moon sets 06:51
  • Sunrise 07:02
  • Moon 98.9%

Night Plan

  • Handover from LBTI to LUCI-AO 00:27
  • OSCO - Run with NCPA 00:30-01:00
  • OT files Location: /home/lbto/Service_Observing/2022A_D_E/LUCI/AO_ESM/
    • ESM_V35.xml
    • AO_Testing_V68.xml
  • On-Axis Full AO (run complete scripts) - 01:00-02:30
    • Bin 3 - Test updates, run with NCPA
      • At least two different stars
    • Bin 1 - Continue to monitor for rolled edges
    • Moon Test - complete before 04hr (Moon too low then)
      • 10 degrees from Moon
        • Bin 1 or Bin 2 10 deg from Moon
          • RA08-09hr folder, Star 13 (BD+21 1788) PA0 -handover-03:15hr
        • Bin 3 or Bin 4 10 deg from Moon
          • RA08-09hr folder, Star 14 (G 51-14) - handover-03:50hr
      • 15 deg from Moon
        • Bin 3 or Bin 4
          • RA08-09hr folder, Star 15 (KUV 08526+3750) PA0 - handover-04:15
  • Bin 2 ESM - confirm issue of defaulting back to Bin 1
    • M67 R=12.2 (change to R=12.6 on script) - visible handover-03:30
    • Dual BLR SDSSJ1713 - R=12.3 (change to R=12.6 on script) - visible 03hr-twilight
  • Bin 1 ESM
    • M61 - face-on spiral R=10.47 star visible handover-twilight
  • Full AO Cat's Eye Nebula - On axis
    • R=11.38 (12.58 - Bin 1 or Bin 2)
    • Visible 03hr-twilight
    • Ks, H2, Pbeta, FeII
  • On-Axis Full AO (run complete scripts) - medium priority
    • Bin 2 - confirm OK
    • Bin 4 - confirm OK
  • Bin 4 ESM - low priority
    • M92 - 02:30-twilight

Poor Weather Backup/Backup in case of AO issues

  • Poor Weather
    • Local Mergers - Rothberg
      • /home/lbto/Service_Observing/2022A_D_E/PoorWeather/Rothberg/LocalMergers/COBandhead_Mergers_V40.xml
      • Any conditions (guiders up to 4" - as long as guider can lock on and WFS) - no particular order
        • UGC 9829 - 2nd half
        • UGC 10607 -2nd half
        • UGC 10675 -2nd half
        • NGC 4004 -2nd half
        • Arp 156 -2nd half
        • NGC 3597 -00:00-03:30hr
    • If clear and super poor-poor seeing
      • LUCI-Bino Standards
        • LUCI - /home/lbto/Service_Observing/2022A_D_E/PoorWeather/LUCI_Phot_Standards/LUCI_Phot_ZP_OT_V3.xml
        • If seeing > 4" - make these track modes in the OT and export
  • If AO cannot be execute and conditions are good
    • LUCI-Bino SL - GRB 211211A - 1.25hr (2nd half) - Rothberg
      • OT File: /home/lbto/Service_Observing/2022A_D_E/LUCI/GRB211211A/GRB211211A_LD2021B-002_V5_OT.xml
      • Ks and J (fraternal) imaging
      • Seeing < 1.3"
      • Thin Clouds
      • If clear (photometric) - execute standard GSPC P272-D before and after observations (just rotates through Ks and J)


Night Log

High winds keeping us closed

10:40 Opening finally. Winds have died down just below the limit, but will have to limit pointing to the north

10:44 Steve pointing binocularly near OSCO target BD+70 622 R7.48

10:50 Seeing 2.3" and we just hit 22.8m/s again. Steve is closing again.

12:17 Opening again. We are the suckers who love a sucker hole. I'm powering up the wavefront sensor while steve opens. He will point and collimate nead BD+70

12:22 Steve is collimating. Seeing is 2.4".....

12:25 Another gust of 22.5m/s, and we are closing. Shutting down the wfs again.

-- BarryRothberg - 15 Feb 2022
Topic revision: r6 - 16 Feb 2022, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback